# Sparfels: Fast Reconstruction from Sparse Unposed Imagery

### v1

@@ Abstract
We present a method for Sparse view reconstruction with surface element splatting that runs within 3 minutes on a consumer grade GPU. While few methods address sparse radiance field learning from noisy or unposed sparse cameras, shape recovery remains relatively underexplored in this setting. Several radiance and shape learning test-time optimization methods address the sparse posed setting by learning data priors or using combinations of external monocular geometry priors. Differently, we propose an efficient and simple pipeline harnessing a single recent 3D foundation model. We leverage its various task heads, notably point maps and camera initializations to instantiate a bundle adjusting 2D Gaussian Splatting (2DGS) model, and image correspondences to guide camera optimization midst 2DGS training. Key to our contribution is a novel formulation of splatted color variance along rays, which can be computed efficiently. Reducing this moment in training leads to more accurate shape reconstructions. We demonstrate state-of-the-art performances in the sparse uncalibrated setting in reconstruction and novel view benchmarks based on established multi-view datasets. Code will be made available at https://shubhendu-jena.github.io/Sparfels/
@@ 1 Introduction
3D reconstruction is a long standing problem at the intersection of computer vision, graphics and machine learning. It is core to a myriad of downstream applications in virtual and augmented reality, autonomous driving, medical imaging, film and animation, to name a few. Classical optimization based techniques such as structure from motion and multi-view stereo [70, 71], shape from visual cues such as shading or silhouette [28, 43], photometric stereo [83] etc., have been popular for decades in this regard. More recently, deep learning has been progressively disrupting 3D vision and top-down approaches to vision more broadly. Among the latest additions, large vision and foundation models (e.g. DUSt3R [79], MASt3R [44]) can offer robustness and fast feed-forward inference compared to classical photogrammetry [70, 71]. Given a few raw uncalibrated images of a scene, they are capable of inferring reasonable relative poses and a competitive coarse 3D structure. These predictions can lack precision and details though, and cannot enable stand-alone novel view synthesis.
@@
By analysing the surface element splatting from a statistical moment perspective, rendered pixels can be seen as color expectation over fragments following the ray termination probability. Based on this, we derive an expression of splatted color variance along the ray, and propose a novel loss tasked with reducing it. This results in improved reconstruction metrics and qualitative sharper mesh details as witnessed by our ablation studies.
The integration of these elements yields state-of-art performance in sparse pose-free 3D reconstruction obtained in very short times on mid range GPUs. We evaluate our method using established reconstruction and novel view datasets [1, 92, 42, 97, 5], showcasing both superior geometric reconstruction, camera pose estimation and novel view synthesis w.r.t. comparable state-of-the-art baselines.
@@ 2 Related work
Traditional NeRFs and GS methods require carefully captured images and SfM preprocessing with COLMAP [69], which could limit practical applications. The few-shot settup has been addressed for both reconstruction from Point Cloud [65, 62, 64, 11] and images. Various approaches address the challenge of limited views through regularization techniques, including depth supervision [17, 58, 76, 46], foundation model supervision [32, 88, 84], specialized architectures [96, 91], while SparseCraft [95] uses learning-free MVS cues for SDF regularization. For Gaussian Splatting, methods like FSGS [104] and SparseGS [86] incorporate external priors. Generalizable methods address sparse-view challenges by pretraining on large datasets, enabling feed-forward inference. Inputs include images (e.g. [10, 12, 49, 45, 35]) and 3D Points (e.g. [9, 52, 60, 30, 63, 61]). For surface reconstruction from color images, approaches like SparseNeuS [50], VolRecon [68], ReTR [47], UfoRecon [56], GeoTransfer [34], and Spurfies [67] leverage various mechanisms such as using image priors to form multi-view cost volumes, using transformer architectures to enhance reconstruction from limited views, as well as using reconstruction priors trained on sparse point clouds.
@@ 3 Method
Given a few unposed color images $\{I^{i}\}_{i=1}^{N}$, our goal is to recover a 3D triangle mesh $\mathcal{S}$ and render novel views of the observed scene efficiently. We achieve this by fitting a bundle adjusting 2DGS [29] model to the images of the scene, bootstrapped by point map and correspondence feedforward predictions from a 3D foundation model (MASt3R [44]) and carefully designed regularization. At test time, mesh $\mathcal{S}$ can be extracted using the TSDF algorithm in Open3D [103] using the converged 2DGS model rendered depth, as recommended in [29]. We detail the workings of our method below.
@@ 3.1 Preliminaries
where $u(\mathbf{p})$ and $v(\mathbf{p})$ are local UV space coordinates of a pixel $\mathbf{p}$ in the primitive’s tangent plane. This ray-splat intersection can be expressed as a funtion of the camera parameters and splat attributes differentiably. Each primitive is parameterized by the center of the Gaussian in world coordinates $\mathbf{x}\in\mathbb{R}^{3}$, a quaternion representing the orientation of the primitive $\mathbf{q}\in\mathbb{R}^{4}$, scale factors for the two dimensions in the local UV tangent plane $\mathbf{s}\in\mathbb{R}^{2}$, opacity $\alpha\in\mathbb{R}$ controlling the transparency of the primitive, and Spherical harmonic coefficients ${\mathbf{c_{\text{sh}}}_{k}\in\mathbb{R}^{3}|k=1,\ldots,n}$ encoding view-dependent appearance. Rotation $\mathbf{q}$ defines a local coordinate frame $[\mathbf{u},\mathbf{v},\mathbf{n}]$, $\mathbf{n}$ being the normal vector orthogonal to the primitive’s plane, and $\mathbf{u}$ and $\mathbf{v}$ spanning the tangent plane.
@@ 3.3 Joint Optimization
We hypothesise that stabilizing and robustifying our training would lead to more robust and improved reconstructions. In order to robustify the training, we would like to control how the rendered color deviates from the ground truth color when the geometry is perturbed, i.e. given divergence measure $\mathcal{D}$, we want to minimize the following worst-case loss along the ray:
@@ 4.1 Datasets and Experimental Configuration
Few-Shot 3D Reconstruction: We evaluate sparse 3D reconstruction on DTU [1], BlendedMVS [92], MVImgNet [97], and MipNeRF360 [5]. For 3-view reconstruction on DTU, we follow SparseNeuS [50], evaluating 15 scenes with two sets of 3-view inputs, reporting the mean per scene. We also provide qualitative comparisons for 3, 6, and 12-view settings on selected scenes from BlendedMVS, MVImgNet, and MipNeRF360.
@@
Camera Pose Estimation: Camera poses are evaluated on Tanks and Temples, MVImgNet, and MipNeRF360, following InstantSplat [21, 22]. Additionally, for few-shot reconstruction, we assess camera poses on DTU using the same 3-view evaluation protocol as in the reconstruction task.
@@
  • NVS specific methods: Pose-free methods NoPe-NeRF [8], CF-3DGS [25], NeRFmm [82], Spar [75] and InstantSplat [21, 22] (in both S and XL configurations). We also include comparisons with 3DGS [40] and FSGS [104], which use pre-computed camera parameters from Colmap [69].
@@
Several key observations can be made from these results. Pretrained generalizable methods (UfoRecon [56] and CasMVSNet [26]) generalize poorly when faced with noisy initial camera parameters from MASt3R, resulting in significantly worse performance. Despite being designed for sparse reconstruction, SparseCraft [95] struggles with the noisy camera poses and point clouds from MASt3R. Spurfies [67], which employs an SDF prior trained on ShapeNet, performs particularly poorly with noisy camera poses, resulting in worse mesh extractions. InstantSplat [21, 22] variants (InstSplatGOF and InstSplat2DGS) perform better than other competing approaches but still produce overly smooth meshes lacking fine details.
Our method significantly outperforms these approaches due to several key advantages, including the variance regularization loss (Sec. 3.3) promoting sharper and more detailed reconstructions Fig 5, The correspondence loss (Sec. 3.3) and joint optimization framework effectively refine the initial noisy camera poses from MASt3R. Our method’s ability to preserve fine geometric details is also evidenced by the superior Normal Consistency scores.
@@
Table 3 presents camera pose estimation results measured by ATE. Our approach achieves significant improvements in camera pose estimation, with a 13.1% reduction in error compared to the best-performing baseline (InstSplatGOF). This demonstrates the effectiveness of our correspondence loss in refining the initial noisy camera poses from MASt3R.
@@ 4.3 Novel View Synthesis Evaluation
BA-2DGS+$\mathcal{L}_{\text{rep}}$+Nomr. init. 7.03 0.829 0.0237
BA-2DGS+$\mathcal{L}_{\text{rep}}$+Nomr. init.+$\mathcal{L}_{\text{var}}$ 6.79 0.834 0.0238
@@ 4.4 Ablation Studies
These experiments were performed using 3 input unposed views on 5 testing scenes from the DTU dataset using two sets of views, following the protocols established in [50, 56]. Tab 4 shows the impact of incorporating different components into our pipeline. Fig 5 shows a qualitative ablation comparison. Camera optimization using bundle adjustment improves all metrics. Incorporating the correspondence loss further enhances this alignment resulting in improved reconstructions. Splat orientation initialization with pointmap normals improves results further. Notice that our variance loss contributes to both the final metrics and the qualitative aspect of our results. The reconstructions with variance loss exhibit significantly sharper details and better preservation of fine structures, validating the quantitative improvements observed in our metrics.
Figure 5: Qualitative ablation of the component of our method, showing also the effect of variance loss. Left blue box: w/o $\mathcal{L}_{\text{var}}$. Right red box: w/o $\mathcal{L}_{\text{var}}$. (3 input images).
@@
We progressively increase the number of views and perform pose-free reconstruction with our method using scan37 from DTU. We evaluate by computing the mean metrics across both test view sets. As shown in Figure 6, our method benefits favourably from increased views. The benefits of all components of our method are also maintained if not increased. This includes our variance loss.
@@ 6 Limitations and Future Work
Figure 7: Qualitative comparison on MVimg and Mipnerf datasets from 6 and 12 input images respectively.
Despite the promising results, several limitations remain, which present opportunities for future research. As foundation models keep on improving, better initializations could be available in the future. Furthermore, while our approach handles a variety of scenes across different datasets, performance may be limited for highly complex geometry, very thin structures, or semi-transparent objects (observed for the glass bottle in Fig 7 room scene), which are common challenges for surface based methods especially in the sparse setting. Additionally, our method assumes a static scene under consistent lighting. Extending it to more dynamic scenarios is one of the challenges we are looking forward to explore.
@@ 7 Conclusion
Sparfels: Fast Reconstruction from Sparse Unposed Imagery
– Supplementary Material –
Shubhendu Jena, Amine Ouasfi, Mae Younes, Adnane Boukhayma
Inria, Univ. Rennes, CNRS, IRISA
Appendix A Evaluation on Novel View Synthesis and Camera pose estimation
In this section, we provide qualitative and quantitative comparisons on the Tanks and Temples [42], MipNeRF360 [5] datasets and MVImgNet [97] datasets for both novel view synthesis and camera pose estimation metrics. Tab 5 presents quantitative results for these experiments on MipNeRF360 and MVImgNet with qualitative results presented in Fig 8 and Fig 9. For Tanks and Temples, quantitative and qualitative results are reported in Tab 6 and Fig 10 respectively. We observe that NoPe-NeRF [8] and NeRF-mm [82] suffer markedly in their novel view performance and camera pose estimation metrics. Being implicit, volumetric rendering methods, they also suffers from slow training and inference times. CF-3DGS [25] also encounters artifacts when rendering from novel viewpoints, stemming from its complex optimization pipeline and erroneous pose estimations. InstantSplat [21, 22] variants provide good performance but still lag behind our method in most metrics, particularly in the challenging 3-view setting. For the Tanks and Temples comparison in Tab 6, Fig 10, we also outperform SPARF [75] by a sizeable margin on all metrics, while requiring order of magnitudes less training and inference time, since it takes around 10 hours to train on a single scene and needs more than a minute to render a single image during inference, owing to it’s volumetric rendering framework. Our method significantly outperforms all baselines on various datasets in terms of SSIM, LPIPS (novel view synthesis metrics) and ATE (camera pose estimation metric), demonstrating its robustness to complex scenes with challenging lighting conditions.
@@ Appendix A Evaluation on Novel View Synthesis and Camera pose estimation
Figure 9: Qualitative comparison of novel view synthesis on MVImgNet [97] dataset.
Figure 10: Qualitative comparison of novel view synthesis on TNT [42] dataset.
@@ Appendix B Additional qualitative comparison on 3D reconstruction
Figure 11: Qualitative comparison with SpaRP [87] on DTU.

### v2

@@ Abstract
We present a method for Sparse view reconstruction with surface element splatting that runs within 3 minutes on a consumer grade GPU. While few methods address sparse radiance field learning from noisy or unposed sparse cameras, shape recovery remains relatively underexplored in this setting. Several radiance and shape learning test-time optimization methods address the sparse posed setting by learning data priors or using combinations of external monocular geometry priors. Differently, we propose an efficient and simple pipeline harnessing a single recent 3D foundation model. We leverage its various task heads, notably point maps and camera initializations to instantiate a bundle adjusting 2D Gaussian Splatting (2DGS) model, and image correspondences to guide camera optimization midst 2DGS training. Key to our contribution is a novel formulation of splatted color variance along rays, which can be computed efficiently. Reducing this moment in training leads to more accurate shape reconstructions. We demonstrate state-of-the-art performances in the sparse uncalibrated setting in reconstruction and novel view benchmarks based on established multi-view datasets. Code will be made available at https://shubhendu-jena.github.io/Sparfels-web/
@@ 1 Introduction
3D reconstruction is a long standing problem at the intersection of computer vision, graphics and machine learning. It is core to a myriad of downstream applications in virtual and augmented reality, autonomous driving, medical imaging, film and animation, to name a few. Classical optimization based techniques such as structure from motion and multi-view stereo [74, 75], shape from visual cues such as shading or silhouette [31, 46], photometric stereo [88] etc., have been popular for decades in this regard. More recently, deep learning has been progressively disrupting 3D vision and top-down approaches to vision more broadly. Among the latest additions, large vision and foundation models (e.g. DUSt3R [84], MASt3R [47]) can offer robustness and fast feed-forward inference compared to classical photogrammetry [74, 75]. Given a few raw uncalibrated images of a scene, they are capable of inferring reasonable relative poses and a competitive coarse 3D structure. These predictions can lack precision and details, though, and cannot enable stand-alone novel view synthesis.
@@
By analyzing the surface element splatting from a statistical moment perspective, rendered pixels can be seen as color expectation over fragments following the ray termination probability. Based on this, we derive an expression of splatted color variance along the ray, and propose a novel loss tasked with reducing it. This results in improved reconstruction metrics and qualitative sharper mesh details as witnessed by our ablation studies.
The integration of these elements yields state-of-art performance in sparse pose-free 3D reconstruction, obtained in very short times on mid-range GPUs. We evaluate our method using established reconstruction and novel view datasets [1, 97, 45, 102, 5], showcasing both superior geometric reconstruction, camera pose estimation and novel view synthesis w.r.t. comparable state-of-the-art baselines.
@@ 2 Related work
Traditional NeRFs and GS methods require carefully captured images and SfM preprocessing with COLMAP [73], which could limit practical applications. The few-shot setup has been addressed for both reconstruction from Point Cloud [69, 66, 68, 12] and images. Various approaches address the challenge of limited views through regularization techniques, including depth supervision [18, 62, 80, 49], foundation model supervision [35, 93, 89], specialized architectures [101, 96], while SparseCraft [100] uses learning-free MVS cues for SDF regularization. For Gaussian Splatting, methods like FSGS [110] and SparseGS [91] incorporate external priors. Generalizable methods address sparse-view challenges by pretraining on large datasets, enabling feed-forward inference. Inputs include images (e.g. [10, 13, 52, 48, 38]) and 3D Points (e.g. [9, 55, 64, 33, 67, 65]). For surface reconstruction from color images, approaches like SparseNeuS [53], VolRecon [72], ReTR [50], UfoRecon [60], GeoTransfer [37], and Spurfies [71] leverage various mechanisms such as using image priors to form multi-view cost volumes, using transformer architectures to enhance reconstruction from limited views, as well as using reconstruction priors trained on sparse point clouds.
@@ 3 Method
Given a few unposed color images $\{I^{i}\}_{i=1}^{N}$, our goal is to recover a 3D triangle mesh $\mathcal{S}$ and render novel views of the observed scene efficiently. We achieve this by fitting a bundle adjusting 2DGS [32] model to the images of the scene, bootstrapped by point map and correspondence feedforward predictions from a 3D foundation model (MASt3R [47]) and carefully designed regularizations. At test time, mesh $\mathcal{S}$ can be extracted using the TSDF algorithm in Open3D [109] using the converged 2DGS model rendered depth, as recommended in [32]. We detail the workings of our method below.
@@ 3.1 Preliminaries
where $u(\mathbf{p})$ and $v(\mathbf{p})$ are local UV space coordinates of a pixel $\mathbf{p}$ in the primitive’s tangent plane. This ray-splat intersection can be expressed as a function of the camera parameters and splat attributes differentiably. Each primitive is parameterized by the center of the Gaussian in world coordinates $\mathbf{x}\in\mathbb{R}^{3}$, a quaternion representing the orientation of the primitive $\mathbf{q}\in\mathbb{R}^{4}$, scale factors for the two dimensions in the local UV tangent plane $\mathbf{s}\in\mathbb{R}^{2}$, opacity $\alpha\in\mathbb{R}$ controlling the transparency of the primitive, and Spherical harmonic coefficients ${\mathbf{c_{\text{sh}}}_{k}\in\mathbb{R}^{3}|k=1,\ldots,n}$ encoding view-dependent appearance. Rotation $\mathbf{q}$ defines a local coordinate frame $[\mathbf{u},\mathbf{v},\mathbf{n}]$, $\mathbf{n}$ being the normal vector orthogonal to the primitive’s plane, and $\mathbf{u}$ and $\mathbf{v}$ spanning the tangent plane.
@@ 3.3 Joint Optimization
We hypothesize that stabilizing and robustifying our training would lead to more robust and improved reconstructions. In order to robustify the training, we would like to control how the rendered color deviates from the ground truth color when the geometry is perturbed, i.e. given divergence measure $\mathcal{D}$, we want to minimize the following worst-case loss along the ray:
@@ 4.1 Datasets and Experimental Configuration
Few-Shot 3D Reconstruction: We evaluate sparse 3D reconstruction on DTU [1], BlendedMVS [97], MVImgNet [102], and MipNeRF360 [5]. For 3-view reconstruction on DTU, we follow SparseNeuS [53], evaluating 15 scenes with two sets of 3-view inputs, reporting the mean per scene. We also provide qualitative comparisons for 3, 6, and 12-view settings on selected scenes from BlendedMVS [97], MVImgNet [102], and MipNeRF360 [5].
@@
Camera Pose Estimation: Camera poses are evaluated on Tanks and Temples [45], MVImgNet [102], and MipNeRF360 [5], following InstantSplat [22, 23]. Additionally, for few-shot reconstruction, we assess camera poses on DTU [1] using the same 3-view evaluation protocol as in the reconstruction task.
@@
  • NVS specific methods: Pose-free methods NoPe-NeRF [8], CF-3DGS [27], NeRFmm [87], SPARF [79] and InstantSplat [22, 23] (in both S and XL configurations). We also include comparisons with 3DGS [43] and FSGS [110], which use pre-computed camera parameters from Colmap [73].
@@
Several key observations can be made from these results. Pretrained generalizable methods (UfoRecon [60] and CasMVSNet [28]) generalize poorly when faced with noisy initial camera parameters from MASt3R [47], resulting in significantly worse performance. Despite being designed for sparse reconstruction, SparseCraft [100] struggles with the noisy camera poses and point clouds from MASt3R [47]. Spurfies [71], which employs an SDF prior trained on ShapeNet, performs particularly poorly with noisy camera poses, resulting in worse mesh extractions. InstantSplat [22, 23] variants (InstSplatGOF and InstSplat2DGS) perform better than other competing approaches, but still produce overly smooth meshes lacking fine details.
Our method significantly outperforms these approaches due to several key advantages, including the variance regularization loss (Sec. 3.3) promoting sharper and more detailed reconstructions Fig 5, The correspondence loss (Sec. 3.3) and joint optimization framework effectively refine the initial noisy camera poses from MASt3R [47]. Our method’s ability to preserve fine geometric details is also evidenced by the superior Normal Consistency scores.
@@
Table 3 presents camera pose estimation results measured by ATE. Our approach achieves significant improvements in camera pose estimation, with a 13.1% reduction in error compared to the best-performing baseline (InstSplatGOF [22]). This demonstrates the effectiveness of our correspondence loss in refining the initial noisy camera poses from MASt3R [47].
@@ 4.3 Novel View Synthesis Evaluation
BA-2DGS+$\mathcal{L}_{\text{rep}}$+Norm. init. 7.03 0.829 0.0237
BA-2DGS+$\mathcal{L}_{\text{rep}}$+Norm. init.+$\mathcal{L}_{\text{var}}$ 6.79 0.834 0.0238
@@ 4.4 Ablation Studies
These experiments were performed using 3 input unposed views on 5 testing scenes from the DTU dataset using two sets of views, following the protocols established in [53, 60]. Tab 4 shows the impact of incorporating different components into our pipeline. Fig 5 shows a qualitative ablation comparison. Camera optimization using bundle adjustment improves all metrics. Incorporating the correspondence loss further enhances this alignment, resulting in improved reconstructions. Splat orientation initialization with pointmap normals improves results further. Notice that our variance loss contributes to both the final metrics and the qualitative aspect of our results. The reconstructions with variance loss exhibit significantly sharper details and better preservation of fine structures, validating the quantitative improvements observed in our metrics.
Figure 5: Qualitative ablation of the component of our method, showing also the effect of variance loss. Left blue box: w/o $\mathcal{L}_{\text{var}}$. Right red box: w $\mathcal{L}_{\text{var}}$. (3 input images).
@@
We progressively increase the number of views and perform pose-free reconstruction with our method using scan37 from DTU. We evaluate by computing the mean metrics across both test view sets. As shown in Figure 6, our method benefits favorably from increased views. The benefits of all components of our method are also maintained, if not increased. This includes our variance loss.
@@ 6 Limitations and Future Work
Figure 7: Qualitative comparison on MVimg and Mipnerf datasets from 6 and 12 input images, respectively.
Despite the promising results, several limitations remain, which present opportunities for future research. As foundation models keep on improving, better initializations could be available in the future. Furthermore, while our approach handles a variety of scenes across different datasets, performance may be limited for highly complex geometry, very thin structures, or semi-transparent objects (observed for the glass bottle in Fig 7 room scene in MipNeRF360 [5]), which are common challenges for surface based methods especially in the sparse setting. Additionally, our method assumes a static scene under consistent lighting. Extending it to more dynamic scenarios is one of the challenges we are looking forward to exploring.
@@ 7 Conclusion
Appendix A Evaluation of Novel View Synthesis and Camera pose estimation
In this section, we provide qualitative and quantitative comparisons on the Tanks and Temples [45], MipNeRF360 [5] datasets and MVImgNet [102] datasets for both novel view synthesis and camera pose estimation metrics. Tab 5 presents quantitative results for these experiments on MipNeRF360 [5] and MVImgNet [102] with qualitative results presented in Fig 8 and Fig 9. For Tanks and Temples [45], quantitative and qualitative results are reported in Tab 6 and Fig 10 respectively. We observe that NoPe-NeRF [8] and NeRF-mm [87] suffer markedly in their novel view performance and camera pose estimation metrics. Being implicit, volumetric rendering methods, they also suffer from slow training and inference times. CF-3DGS [27] also encounters artifacts when rendering from novel viewpoints, stemming from its complex optimization pipeline and erroneous pose estimations. InstantSplat [22, 23] variants provide good performance, but still lag behind our method in most metrics, particularly in the challenging 3-view setting. For the Tanks and Temples comparison in Tab 6, Fig 10, we also outperform SPARF [79] by a sizeable margin on all metrics, while requiring order of magnitudes less training and inference time, since it takes around 10 hours to train on a single scene and needs more than a minute to render a single image during inference, owing to its volumetric rendering framework. Our method significantly outperforms all baselines on various datasets in terms of SSIM, LPIPS (novel view synthesis metrics) and ATE (camera pose estimation metric), demonstrating its robustness to complex scenes with challenging lighting conditions.
@@ Appendix A Evaluation of Novel View Synthesis and Camera pose estimation
Figure 9: Qualitative comparison of novel view synthesis on MVImgNet [102] dataset from 3 input images.
Figure 10: Qualitative comparison of novel view synthesis on TNT [45] dataset from 3 input images.
@@ Appendix B Additional qualitative comparison on 3D reconstruction
Figure 11: Qualitative comparison with SpaRP [92] on DTU from 3 input images.
Appendix C Color variance plot
We plot the average color variance over optimization iterations (Fig. 12) for models w/ and w/o variance loss. Models with the loss activated effectively maintain lower color variance consistently, which aligns with our goal of encouraging stable, low-uncertainty renderings. This supports the effectiveness of the proposed loss in guiding convergence toward robust geometry.
Figure 12: Color variance. The variance loss keeps color variance lower, encouraging stable, robust convergence.
Appendix D Alternative priors
This example (Fig. 13) demonstrates that initializing our framework with VGGT [81], a recent state-of-the-art feed-forward method that avoids the global optimization step of MASt3R [47] also produces successful results. This highlights the modularity of our approach and its compatibility with different geometric priors.
Figure 13: Alternative prior. VGGT initialization shows successful results (3 input images), demonstrating our framework’s modularity.
Appendix E Variance loss motivation
Our goal is to hedge against epistemic uncertainty in geometry estimation inherent to the unposed surface reconstruction problem. Under sparse-view supervision, the rendering objective admits many geometries that fit the training images but generalizes poorly (see Sec.3 in [107]). This issue is exacerbated when camera poses are optimized during training, introducing additional noise into the supervision. In Gaussian Splatting, scene geometry is encoded through splat parameters defining the 3D density. Among the many plausible geometries, we seek to bias the model toward those that remain predictive even under small perturbations, i.e. robust solutions less sensitive to noisy supervision signals. To formalize this, we minimize the worst-case deviation in rendered color under perturbations to the geometric density field (Eq. 8), yielding a variance regularization loss (Eq. 10) that penalizes color variance along rays. From a learning-theoretic perspective, this can be interpreted as seeking flat minima [29] in the space of densities, an idea supported by arguments from both statistical and deep learning viewpoints, and shown to be effective across a range of machine learning applications [11, 56, 29, 26]. Empirically, this leads to more stable and consistent reconstructions from sparse views (Fig. 5, Tab. 4).
Appendix F Prior failures
The example below illustrates a typical scenario where MASt3R’s [47] feed-forward geometry prediction struggles: reconstruction from only 6 images without known camera poses. Challenging regions such as texture-less surfaces, highly reflective materials (e.g., glass doors and shiny faucets), and thin structures often lead to noisy or incomplete results. In contrast, our method recovers plausible geometry in these cases thanks to robust test-time optimization, which refines both the pose and the reconstructed shape despite imperfect initializations.
Figure 14: Robustness to MASt3R failure (3 input images). Our method recovers geometry where MASt3R struggles.